OpenStudy (anonymous):

if a ball is thrown vertically upward with a velocity of 80ft/s, then its height after t seconds is s=80t-26t^2. how do i find the maximum height reached by the ball?

OpenStudy (anonymous):

For some reason I was thinking s was velocity. :-/

OpenStudy (anonymous):

OOps, wrong! x = 80/ 52 = 1.5 ft

OpenStudy (anonymous):

@Chlorophyll what's your method?

OpenStudy (anonymous):

And what does s represent? Distance?

OpenStudy (anonymous):

maximum of parabola has vertex ( x = -b/2a, y = f ( -b/2a ) )

OpenStudy (anonymous):

oh okay... for some reason I was thinking about kinematics equations.

OpenStudy (anonymous):

Yes, "its heigt = s..."

OpenStudy (anonymous):

take the first derivative of s (80-52t) find the zero (1.5384615) that is the time at which te height is reaced then plug that into the original equation its 61.53 ft

OpenStudy (anonymous):

Derivative's used for calculus level, while vertex applies for algebra level!

OpenStudy (anonymous):

It's so tempting to use calc for these problems

OpenStudy (anonymous):

it gives the same answer im just usingthe esiest way for me to do you could always jsut graph the quadratic and find the maximum either way is the same

OpenStudy (earthcitizen):

kinematics

OpenStudy (anonymous):

But then I remember all those kinematics equations from physics and keep thinking I need to use those (banging head on desk right now).

OpenStudy (anonymous):

Yep, but I'm unsure if the asker at that level .

OpenStudy (anonymous):

It's gotta be algebra level. I remembered having to do probs like these 6 yrs ago when I took it.

OpenStudy (anonymous):

I'm horrible at Physics, teach me some day, will you, people !

OpenStudy (anonymous):

no matter how you do it its 61.53846 to 7 sig figs lol

OpenStudy (anonymous):

Why don't we just find the vertex of the parabola. \[{\rm \max~height} = {-b \over 2a}\]

OpenStudy (anonymous):

80/52

OpenStudy (anonymous):

that solves for x and the question wants y so you still have to plug that into the original equation to find theheight not the time that that height occurs

OpenStudy (anonymous):

Then plug it back in and solve for s.

OpenStudy (anonymous):

it's calculus and i noticed i wrote the equation wrong its s=80t-16t^2 and i think i figured it out, i first took the derivative of it which was v(t)= 80-32t then solved for 0 because the velocity at max is zero then i plugged what i found for t into the original equation which was 2.5 so it was 80(2.5)-16(2.5)^2 and it ends up to be 100ft

OpenStudy (anonymous):

that sthe correctway to do the problem, and yes changing 26 to 16 will change the answer LOL

OpenStudy (anonymous):

for this same equation, i need to find the velocity of the ball when its at 96 ft above the ground on its way up and its way down?

OpenStudy (earthcitizen):

differentiate the equation again to find the acceleration, then use v=u+at